If a medication has a concentration of 5mg per 1.5ml, then how many ml is 0.3mg?
Answer by CubeyThePenguin(3113)

5 mg/1.5 ml = 0.3 mg/x ml

5/1.5 = 0.3/x

x = (1.5 * 0.3)/5 = 0.09 mL